• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Punjabi

    Log entry time 00:06:51 on June 5,2002

    Entry number 85255

    keyword=Sift Summary

    16:00 Beam is off due to maintenance.

    20:40 beam back. Asked for 15 microamps.

    20:45 Start run 1164 in Q1_Acc kinematics. Both collimators open.

    21:05 start run 1165. Q1-Acc kinematics. 6 msr left collimator in. Right arm open collimator.

    Target charts for ESR temperature etc. were fozen (value not changing). Doug Rebooted target IOC. Some problems for high power heater to restart in both loops. Doug called Kathy and fixed things. See separate entry from Doug.

    21:40 Started run 1166. Q1_Acc kinematics. Both spectrometers now have 6 msr collimator in place.

    22:00 stopped run 1166.

    Asked MCC to stop raster and go to 5 microamp. Changed target to optics to do harp measurement.

    Harp run 1768, CODA run 1167
    Harp run 1769 CODA run 1167.

    Harp measurements: Horizontal sigma 112 microm
    Vertical sigma 59.64 microm.

    22:18 Stopped harp measurement.

    22:18 Changing target to 4 cm hydrogen.

    22:32 Start run 1168. Q1_Acc kinematics. 4 cm hydrogen target. Both spectrometer with 6 msr collimator in place.

    22:51 stopped run 1168.

    22:54 Start run 1169. Q1_Acc kinematics. Right spectrometer collimator open. Left spectrometer collimator 6 msr.

    23:10 stopped run 1169.

    23:13 Started run 1170. Q1_Acc kinematics. Both spectrometers collimator open.

    23:29 stopped run 1170.

    23:37 start run 1171. Q1_Acc kinematics. Left spectrometer sieve slit in. Right spectrometer open collimator.

    23:59 stopped run 1171. Lost beam.